OpenStudy (anonymous):

determine if series is absolutely convergent. sum from 1 to infinity (-6)^n/n! Here's what I did so far lim as n approaches infinity (-6)^(n+1)/ (n+1)! x n!/(-6)^n I did some canceling and got limit as n approaches infinity absolute value of -6n!/ (n+1)! When i plug in n for infinity i got a indeterminate form. Answer suppose to be absolutely convergent. what did i do wrong. Help!

OpenStudy (anonymous):

Cancel out the factorials: (n+1)!=(n+1)n!

OpenStudy (anonymous):

where did you get the extra n?

OpenStudy (anonymous):

The definition of the factorial is n! = n * (n-1) * (n-2) ... 3*2*1 The product has to have exactly n factors. This means that the factorial of n+1 has one more factor in it: (n+1)! = (n+1) * n * (n-1) * (n-2) ... 3*2*1 Since the last n factors are what we call n!, we can write (n+1)! = (n+1) * n!